# On $[1,2]$-Domination in Interval and Circle Graphs

### v2

@@ 2 Algorithm for interval graphs
Conversely, let $j\in[\mathtt{maxlow}(i),i^{\prime})$. Assume that $S$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S$ and $a\notin S$ for all $a\in[\mathtt{low}(j+1,i),j)$. So, $|S|=\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j)$. Since $j\in[\mathtt{maxlow}(i),i^{\prime})$ and $a\notin S$ for all $a\in[\mathtt{low}(j+1,i),j)$, each vertex of $[i^{\prime},i]$ is exactly dominated by $j$. We obtain that $S$ is an $[1,2]$-dominating set of $G[1,i]$ such that $a\notin S$ for all $a\in[i^{\prime},i]$, that is, $|S|\leq\gamma^{0}_{[1,2]}(i^{\prime},i)$. Let $k\in[\mathtt{low}(j+1,i),j)$. Assume that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j,k\in S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[\mathtt{low}(j+1,i),j)\setminus\{k\}$. So, $|S^{\prime}|=\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j,k)$. Since $j\in[\mathtt{maxlow}(i),i^{\prime})$, $k\in[\mathtt{low}(j+1,i),j)\cap S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[\mathtt{low}(j+1,i),j)\setminus\{k\}$, each vertex of $[i^{\prime},i]$ is dominated by at least one and at most two vertices of $S^{\prime}$. We obtain that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $a\notin S^{\prime}$ for all $a\in[i^{\prime},i]$, that is, $|S^{\prime}|\leq\gamma^{0}_{[1,2]}(i^{\prime},i)$. Hence, $\min\{\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j)=|S|,\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j,k)=|S^{\prime}|:j\in[\mathtt{maxlow}(i),i^{\prime}),k\in[\mathtt{low}(j+1,i),j)\}\leq\gamma^{0}_{[1,2]}(i^{\prime},i)$. This completes the proofs of the case (ii) and the lemma. ∎
Lemma 3.
@@
Proof.
Let $D$ be an $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality such that $k\in D$ and $a\notin D$ for all $a\in[j,i]\setminus\{k\}$. So, $|D|=\gamma^{0}_{[1,2]}(j,i:k)$. Since $a\notin D$ for all $a\in(k,i]$, it obtains that $D$ is an $[1,2]$-dominating set of $G[1,k]$. By the case (ii) of Corollary 2, there is a vertex $x\in[\mathtt{maxlow}(i),i]$ such that $xy\notin E$ for each vertex $y\in[1,\mathtt{maxlow}(i))$. Thus, if $k<\mathtt{maxlow}(i)$, then $D$ is not an $[1,2]$-dominating set of $G[1,i]$, that is, $\gamma^{0}_{[1,2]}(j,i:k)$ is not defined. Assume that $\mathtt{maxlow}(i)\leq k$. Let $c\in(k,i]$ be a vertex such that $\mathtt{low}(c)\leq\mathtt{low}(a)$ for all $a\in(k,i]$, that is, $\mathtt{low}(c)=\mathtt{low}(k+1,i)$. See Figure 6. Recall that $b=\min\{\mathtt{low}(k+1,i),j\}$ and $a\notin D$ for all $a\in[j,i]\setminus\{k\}$. If $\mathtt{low}(k+1,i)<j$, that is $b=\mathtt{low}(k+1,i)$, then because $c\notin D$ and $k\in D$, at most one of vertices of $[b,j)$ is in $D$, otherwise, $b=j$ and no vertex of $[b,k)$ is in $D$. Hence, either no vertex of $[b,k)$ is in $D$, see Figure 6(a), or exactly one vertex of $[b,j)$ is in $D$ (if $b<j$), see Figure 6(b). In the following we consider these two cases.
@@
Let $l\in[b,j)$. Assume that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,k]$ with minimum cardinality such that $l,k\in S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[b,k)\setminus\{l\}$. So, $|S^{\prime}|=\gamma^{1}_{[1,2]}(b,k:k,l)$. Since $l\in[b,j)$, $k\in[\mathtt{maxlow}(i),i)$ and $a\notin S^{\prime}$ for all $a\in[b,i)\setminus\{k,l\}$, each vertex of $(k,i]$ is dominated by at least one and at most two vertices of $S^{\prime}$. We obtain that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $k\in S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[j,i]\setminus\{k\}$, that is, $|S^{\prime}|\leq\gamma^{0}_{[1,2]}(j,i:k)$. Hence, $\gamma^{0}_{[1,2]}(j,i:k)=\min\{\gamma^{1}_{[1,2]}(b,k:k)=|S|,\gamma^{1}_{[1,2]}(b,k:k,l)=|S^{\prime}|:l\in[b,j)\}$ This completes the proof of the lemma. ∎
Lemma 4.
@@
Proof.
Suppose that $j=max[\mathtt{low}(i),i^{\prime\prime})\setminus D$. See Figure 7. Since $j\notin D$ and $j$ is adjacent to $i$, the vertex $j$ is adjacent to at most one vertex of $S=D\cap(j,i)$. If $j$ is adjacent to no vertex of $S$, then $D^{\prime}=D\setminus S$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i\in D^{\prime}$, a contradiction. Let $j$ be adjacent to exactly one vertex $k\in S$. We obtain that $D^{\prime\prime}=D\setminus S\cup\{k\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i\in D^{\prime}$, a contradiction. This completes the proof of the lemma. ∎
@@
Assume that at least three vertices of $[b,i)$ are in $D$ and let $i^{\prime}=\max D\cap[1,i)$ and $i^{\prime\prime}=\max D\cap[1,i^{\prime})$. By Lemma 4, $x\in D$ for all $x\in[b,i^{\prime\prime})$. Hence,
@@
Clearly, $\gamma^{1}_{[1,2]}(i)\leq\min\{\gamma^{1}_{[1,2]}(b,i:i),\gamma^{1}_{[1,2]}(b,i:i,j),\gamma^{1}_{[1,2]}(b,i:i,j^{\prime},k^{\prime}),\gamma^{11}_{[1,2]}(b,i:i,j^{\prime},k^{\prime}):b\leq j<i,b\leq k^{\prime}<j^{\prime}<i\}$. This, together with inequalities (1)–(4), completes the proof of the lemma. ∎
Lemma 6.
@@
Proof.
Let $D$ be an $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality such that $i\in D$ and $x\notin D$ for all $x\in[j,i)$. So, $|D|=\gamma^{1}_{[1,2]}(j,i:i)$. If $j\leq\mathtt{low}(b-1)$, then there is no vertex in $D$ to dominate $b-1$ and so $D$ is not an $[1,2]$-dominating set of $G[1,i]$, that is, $\gamma^{1}_{[1,2]}(j,i:i)$ is not defined. See Figure 8(a). This proves the case (i). In the rest of the proof, we assume that $\mathtt{low}(b-1)<j$. Recall that $a=\max\{j,b\}$. Let $c\in[a,i-1)$ be a vertex such that $\mathtt{low}(c)\leq\mathtt{low}(x)$ for all $x\in[a,i-1)$, that is, $\mathtt{low}(c)=\mathtt{low}(a,i-1)$. We distinguish two cases depending on $j<\mathtt{low}(c)$ or $\mathtt{low}(c)\leq j$. In the following we consider these cases.
@@
    Conversely, assume that $S$ is an $[1,2]$-dominating set of $G[1,b-1]$ with minimum cardinality such that $x\notin S$ for all $x\in[j,b-1]$. So, $|S|=\gamma^{0}_{[1,2]}(j,b-1)$. We see that $S^{\prime}=S\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i\in S^{\prime}$ and $x\notin D$ for all $x\in[j,i-1]$. So, $\gamma^{1}_{[1,2]}(j,i:i)\leq|S^{\prime}|=\gamma^{0}_{[1,2]}(j,b-1)+1$. This completes the proof of the case (ii).
@@
      Assume that $x\notin D$ for all $x\in[\mathtt{low}(c),j-1]$. We obtain that $D^{\prime}=D\setminus\{i\}$ is an $[1,2]$-dominating set of $G[1,j-1]$ such that $x\notin D^{\prime}$ for all $x\in[\mathtt{low}(c),j-1]$. Hence,
@@
      Assume that exactly one vertex $k\in[\mathtt{low}(c),j-1]$ is in $D$. If $k<j-1$, then $D^{\prime}=D\setminus\{i\}$ is an $[1,2]$-dominating set of $G[1,j-1]$ such that $k\in D^{\prime}$ and $x\notin D^{\prime}$ for all $x\in[\mathtt{low}(c),j-1]$. Hence,
@@
      If $k=j-1$, then $D^{\prime}=D\setminus\{i\}$ is an $[1,2]$-dominating set of $G[1,j-1]$ such that $j-1\in D^{\prime}$ and $x\notin D^{\prime}$ for all $x\in[\mathtt{low}(c),j-1]$. Hence,
@@
This completes the proof of the lemma. ∎
Lemma 7.
@@
Recall that $d=\mathtt{low}(j+1,i-1)$ and $z=\max\{x\in(j,i):d=\mathtt{low}(x)\}$. If $j\geq b$, then $z\geq b$. Hence, we have (I) $j<b$ and $z<b$, (II) $j<b$ and $z\geq b$, or (III) $j\geq b$. In the following we consider these cases.
@@
    Assume $k=j\leq d$. See Figure 9(b). We get that $D^{\prime}=D\setminus\{i\}$ is an $[1,2]$-dominating set of $G[1,j]$ such that $j\in D^{\prime}$. Hence, $\gamma^{1}_{[1,2]}(j)\leq|D^{\prime}|=\gamma^{1}_{[1,2]}(k,i:i,j)-1$. Conversely, assume that $S$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S$. So, $|S|=\gamma^{1}_{[1,2]}(j)$. Since $j\leq d$ and the case (i) does not hold, each vertex of $(j,i)$ is adjacent to at most two vertices $i$ and $j$. Hence, $S^{\prime}=S\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}$ and $x\notin S^{\prime}$ for all $x\in[k,i)\setminus\{j\}$ and so $\gamma^{1}_{[1,2]}(k,i:i,j)\leq|S^{\prime}|=\gamma^{1}_{[1,2]}(j)+1$. This completes the proof of the case (ii.a).
@@
    Assume $k<j$. See Figure 9(c). Clearly, $y$ is adjacent to $i$. We get that $D^{\prime}=D\setminus\{i\}$ is an $[1,2]$-dominating set of $G[1,j]$ such that $j\in D^{\prime}$ and $x\notin D^{\prime}$ for all $x\in[\min\{k,d^{\prime\prime}\},j)$. Hence, $\gamma^{1}_{[1,2]}(\min\{k,d^{\prime\prime}\},j:j)\leq|D^{\prime}|=\gamma^{1}_{[1,2]}(k,i:i,j)-1$. Conversely, assume that $S$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S$ and $x\notin S$ for all $x\in[\min\{k,d^{\prime\prime}\},j)$. So, $|S|=\gamma^{1}_{[1,2]}(\min\{k,d^{\prime\prime}\},j:j)$. Since $x\notin S$ for all $x\in[\min\{k,d^{\prime\prime}\},j)$, $k\leq d$ and the case (i) does not hold, each vertex of $(j,i)$ is adjacent to at most two vertices $i$ and $j$. Hence, $S^{\prime}=S\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}$ and $x\notin S^{\prime}$ for all $x\in[\min\{k,d^{\prime\prime}\},i)\setminus\{j\}$ and so $\gamma^{1}_{[1,2]}(k,i:i,j)\leq|S^{\prime}|=\gamma^{1}_{[1,2]}(\min\{k,d^{\prime\prime}\},j:j)+1$. This completes the proof of the case (ii.b).
@@
    If no vertex of $[d,\min\{k,d^{\prime\prime}\})$ is in $D$, then we get that $D^{\prime}_{1}=D\setminus\{i\}$ is an $[1,2]$-dominating set of $G[1,j]$ such that $j\in D^{\prime}_{1}$ and $x\notin D^{\prime}_{1}$ for all $x\in[d,j)$. Hence,
@@
    If exactly one vertex $w\in[d,\min\{k,d^{\prime\prime}\})$ is in $D$, then we get that $D^{\prime}_{2}=D\setminus\{i\}$ is an $[1,2]$-dominating set of $G[1,j]$ such that $j,w\in D^{\prime}_{2}$ and $x\notin D^{\prime}_{2}$ for all $x\in[d,j)\setminus\{w\}$. Assume that $x^{\prime}\in[d,\min\{k,d^{\prime\prime}\})$. Hence,
@@
    Conversely, assume that $S_{1}$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S_{1}$ and $x\notin S_{1}$ for all $x\in[d,j)$. So, $|S_{1}|=\gamma^{1}_{[1,2]}(d,j:j)$. Since $x\notin S_{1}$ for all $x\in[d,j)$, $d<k$ and the case (i) does not hold, each vertex of $(j,i)$ is adjacent to at most two vertices $i$ and $j$. Hence, $S^{\prime}_{1}=S_{1}\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}_{1}$ and $x\notin S^{\prime}_{1}$ for all $x\in[k,i)\setminus\{j\}$ and so
@@
    Let $x\in[d,\min\{k,d^{\prime\prime}\})$. Assume that $S_{2}$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j,x\in S_{2}$ and $w\notin S_{2}$ for all $w\in[d,j)\setminus\{x\}$. So, $|S_{2}|=\gamma^{1}_{[1,2]}(d,j:j,x)$. We get that $S^{\prime}_{2}=S_{2}\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}_{2}$ and $w\notin S^{\prime}_{2}$ for all $w\in[k,i)\setminus\{j\}$ and so
@@
Case II. Assume that $j<b$ and $z\geq b$. The proofs of the cases (iii.a) and (iii.c) is similar to the proof of the case (ii.a) and the proof of the case (iii.b) is similar to the proof of the case (ii.b), respectively.
@@
    Assume that $c<b$. Recall that $k\leq j$, $d=\mathtt{low}(j+1,i-1)$, $d^{\prime}=\mathtt{low}(b,j-1)$ and $k^{\prime}=\min\{k,d,d^{\prime}\}$. Clearly, $k^{\prime}$ exists. Since $b\leq j$, $k^{\prime}\leq b$. We claim that all vertices of $[k^{\prime},j)$ are not in $D$. If $k^{\prime}=k$, the claim is clear. Assume that $k^{\prime}=d$. So, $d\leq b$. Recall that $z=\max\{x\in(j,i):d=\mathtt{low}(x)\}$. We obtain that $z\notin D$ is adjacent to $i,j\in D$. Since $D$ is an $[1,2]$-dominating set, all vertices of $[d,j)$ are not in $D$. Assume that $k^{\prime}=d^{\prime}$. Let $z^{\prime}\in[b,j-1]$ such that $\mathtt{low}(z^{\prime})=d^{\prime}$. Since $c<b$, we get that $z^{\prime}\notin D$ is adjacent to $i,j\in D$ and so all vertices of $[d^{\prime},j)$ are not in $D$. Hence, all vertices of $[k^{\prime},j)$ are not in $D$. The proof of the case (iv.c) is similar to the proof of the case (ii.b).
This completes the proof of the lemma. ∎

### v4

@@ 2 Algorithm for interval graphs
Conversely, let $j\in[\mathtt{maxlow}(i),i^{\prime})$. Assume that $S$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S$ and $a\notin S$ for all $a\in[\mathtt{low}(j+1,i),j)$. So, $|S|=\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j)$. Since $j\in[\mathtt{maxlow}(i),i^{\prime})$ and $a\notin S$ for all $a\in[\mathtt{low}(j+1,i),j)$, each vertex of $[i^{\prime},i]$ is exactly dominated by $j$. We obtain that $S$ is an $[1,2]$-dominating set of $G[1,i]$ such that $a\notin S$ for all $a\in[i^{\prime},i]$, that is, $|S|\leq\gamma^{0}_{[1,2]}(i^{\prime},i)$. Let $k\in[\mathtt{low}(j+1,i),j)$. Assume that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j,k\in S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[\mathtt{low}(j+1,i),j)\setminus\{k\}$. So, $|S^{\prime}|=\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j,k)$. Since $j\in[\mathtt{maxlow}(i),i^{\prime})$, $k\in[\mathtt{low}(j+1,i),j)\cap S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[\mathtt{low}(j+1,i),j)\setminus\{k\}$, each vertex of $[i^{\prime},i]$ is dominated by at least one and at most two vertices of $S^{\prime}$. We obtain that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $a\notin S^{\prime}$ for all $a\in[i^{\prime},i]$, that is, $|S^{\prime}|\leq\gamma^{0}_{[1,2]}(i^{\prime},i)$. Hence, $\min\{\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j)=|S|,\gamma^{1}_{[1,2]}(\mathtt{low}(j+1,i),j:j,k)=|S^{\prime}|:j\in[\mathtt{maxlow}(i),i^{\prime}),k\in[\mathtt{low}(j+1,i),j)\}\leq\gamma^{0}_{[1,2]}(i^{\prime},i)$. This completes the proofs of the case (ii) and the lemma.
Lemma 2.2.
@@
Proof 2.3.
Let $D$ be a $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality such that $k\in D$ and $a\notin D$ for all $a\in[j,i]\setminus\{k\}$. So, $|D|=\gamma^{0}_{[1,2]}(j,i:k)$. Since $a\notin D$ for all $a\in(k,i]$, it obtains that $D$ is an $[1,2]$-dominating set of $G[1,k]$. By the case (ii) of Corollary 3, there is a vertex $x\in[\mathtt{maxlow}(i),i]$ such that $xy\notin E$ for each vertex $y\in[1,\mathtt{maxlow}(i))$. Thus, if $k<\mathtt{maxlow}(i)$, then $D$ is not an $[1,2]$-dominating set of $G[1,i]$, that is, $\gamma^{0}_{[1,2]}(j,i:k)$ is not defined. Assume that $\mathtt{maxlow}(i)\leq k$. Let $c\in(k,i]$ be a vertex such that $\mathtt{low}(c)\leq\mathtt{low}(a)$ for all $a\in(k,i]$, that is, $\mathtt{low}(c)=\mathtt{low}(k+1,i)$. See Figure 6. Recall that $b=\min\{\mathtt{low}(k+1,i),j\}$ and $a\notin D$ for all $a\in[j,i]\setminus\{k\}$. If $\mathtt{low}(k+1,i)<j$, that is $b=\mathtt{low}(k+1,i)$, then because $c\notin D$ and $k\in D$, at most one of vertices of $[b,j)$ is in $D$, otherwise, $b=j$ and no vertex of $[b,k)$ is in $D$. Hence, either no vertex of $[b,k)$ is in $D$, see Figure 6(a), or exactly one vertex of $[b,j)$ is in $D$ (if $b<j$), see Figure 6(b). In the following we consider these two cases.
@@
Let $l\in[b,j)$. Assume that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,k]$ with minimum cardinality such that $l,k\in S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[b,k)\setminus\{l\}$. So, $|S^{\prime}|=\gamma^{1}_{[1,2]}(b,k:k,l)$. Since $l\in[b,j)$, $k\in[\mathtt{maxlow}(i),i)$ and $a\notin S^{\prime}$ for all $a\in[b,i)\setminus\{k,l\}$, each vertex of $(k,i]$ is dominated by at least one and at most two vertices of $S^{\prime}$. We obtain that $S^{\prime}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $k\in S^{\prime}$ and $a\notin S^{\prime}$ for all $a\in[j,i]\setminus\{k\}$, that is, $|S^{\prime}|\leq\gamma^{0}_{[1,2]}(j,i:k)$. Hence, $\gamma^{0}_{[1,2]}(j,i:k)=\min\{\gamma^{1}_{[1,2]}(b,k:k)=|S|,\gamma^{1}_{[1,2]}(b,k:k,l)=|S^{\prime}|:l\in[b,j)\}$ This completes the proof of the lemma.
Lemma 2.4.
@@
Proof 2.5.
Suppose that $j=max[\mathtt{low}(i),i^{\prime\prime})\setminus D$. See Figure 7. Since $j\notin D$ and $j$ is adjacent to $i$, the vertex $j$ is adjacent to at most one vertex of $S=D\cap(j,i)$. If $j$ is adjacent to no vertex of $S$, then $D^{\prime}=D\setminus S$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i\in D^{\prime}$, a contradiction. Let $j$ be adjacent to exactly one vertex $k\in S$. We obtain that $D^{\prime\prime}=D\setminus S\cup\{k\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i\in D^{\prime}$, a contradiction. This completes the proof of the lemma.
@@
Assume that at least three vertices of $[b,i)$ are in $D$ and let $i^{\prime}=\max D\cap[1,i)$ and $i^{\prime\prime}=\max D\cap[1,i^{\prime})$. By Lemma 2.4, $x\in D$ for all $x\in[b,i^{\prime\prime})$. Hence,
@@
Clearly, $\gamma^{1}_{[1,2]}(i)\leq\min\{\gamma^{1}_{[1,2]}(b,i:i),\gamma^{1}_{[1,2]}(b,i:i,j),\gamma^{1}_{[1,2]}(b,i:i,j^{\prime},k^{\prime}),\gamma^{11}_{[1,2]}(b,i:i,j^{\prime},k^{\prime}):b\leq j<i,b\leq k^{\prime}<j^{\prime}<i\}$. This, together with inequalities (1)–(4), completes the proof of the lemma.
Lemma 2.8.
@@
Proof 2.9.
Let $D$ be an $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality such that $i\in D$ and $x\notin D$ for all $x\in[j,i)$. So, $|D|=\gamma^{1}_{[1,2]}(j,i:i)$. If $j\leq\mathtt{low}(b-1)$, then there is no vertex in $D$ to dominate $b-1$ and so $D$ is not a $[1,2]$-dominating set of $G[1,i]$, that is, $\gamma^{1}_{[1,2]}(j,i:i)$ is not defined. See Figure 8(a). This proves the case (i). In the rest of the proof, we assume that $\mathtt{low}(b-1)<j$. Recall that $a=\max\{j,b\}$. Let $c\in[a,i-1)$ be a vertex such that $\mathtt{low}(c)\leq\mathtt{low}(x)$ for all $x\in[a,i-1)$, that is, $\mathtt{low}(c)=\mathtt{low}(a,i-1)$. We distinguish two cases depending on $j<\mathtt{low}(c)$ or $\mathtt{low}(c)\leq j$. In the following we consider these cases.
@@
    Conversely, assume that $S$ is a $[1,2]$-dominating set of $G[1,b-1]$ with minimum cardinality such that $x\notin S$ for all $x\in[j,b-1]$. So, $|S|=\gamma^{0}_{[1,2]}(j,b-1)$. We see that $S^{\prime}=S\cup\{i\}$ is a $[1,2]$-dominating set of $G[1,i]$ such that $i\in S^{\prime}$ and $x\notin D$ for all $x\in[j,i-1]$. So, $\gamma^{1}_{[1,2]}(j,i:i)\leq|S^{\prime}|=\gamma^{0}_{[1,2]}(j,b-1)+1$. This completes the proof of the case (ii).
@@
      Assume that $x\notin D$ for all $x\in[\mathtt{low}(c),j-1]$. We obtain that $D^{\prime}=D\setminus\{i\}$ is a $[1,2]$-dominating set of $G[1,j-1]$ such that $x\notin D^{\prime}$ for all $x\in[\mathtt{low}(c),j-1]$. Hence,
@@
      Assume that exactly one vertex $k\in[\mathtt{low}(c),j-1]$ is in $D$. If $k<j-1$, then $D^{\prime}=D\setminus\{i\}$ is a $[1,2]$-dominating set of $G[1,j-1]$ such that $k\in D^{\prime}$ and $x\notin D^{\prime}$ for all $x\in[\mathtt{low}(c),j-1]$. Hence,
@@
      If $k=j-1$, then $D^{\prime}=D\setminus\{i\}$ is a $[1,2]$-dominating set of $G[1,j-1]$ such that $j-1\in D^{\prime}$ and $x\notin D^{\prime}$ for all $x\in[\mathtt{low}(c),j-1]$. Hence,
@@
This completes the proof of the lemma.
Lemma 2.10.
@@
Recall that $d=\mathtt{low}(j+1,i-1)$ and $z=\max\{x\in(j,i):d=\mathtt{low}(x)\}$. If $j\geq b$, then $z\geq b$. Hence, we have (I) $j<b$ and $z<b$, (II) $j<b$ and $z\geq b$, or (III) $j\geq b$. In the following, we consider these cases.
@@
    Assume $k=j\leq d$. See Figure 9(b). We get that $D^{\prime}=D\setminus\{i\}$ is a $[1,2]$-dominating set of $G[1,j]$ such that $j\in D^{\prime}$. Hence, $\gamma^{1}_{[1,2]}(j)\leq|D^{\prime}|=\gamma^{1}_{[1,2]}(k,i:i,j)-1$. Conversely, assume that $S$ is a $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S$. So, $|S|=\gamma^{1}_{[1,2]}(j)$. Since $j\leq d$ and the case (i) does not hold, each vertex of $(j,i)$ is adjacent to at most two vertices $i$ and $j$. Hence, $S^{\prime}=S\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}$ and $x\notin S^{\prime}$ for all $x\in[k,i)\setminus\{j\}$ and so $\gamma^{1}_{[1,2]}(k,i:i,j)\leq|S^{\prime}|=\gamma^{1}_{[1,2]}(j)+1$. This completes the proof of the case (ii.a).
@@
    Assume $k<j$. See Figure 9(c). Clearly, $y$ is adjacent to $i$. We get that $D^{\prime}=D\setminus\{i\}$ is a $[1,2]$-dominating set of $G[1,j]$ such that $j\in D^{\prime}$ and $x\notin D^{\prime}$ for all $x\in[\min\{k,d^{\prime\prime}\},j)$. Hence, $\gamma^{1}_{[1,2]}(\min\{k,d^{\prime\prime}\},j:j)\leq|D^{\prime}|=\gamma^{1}_{[1,2]}(k,i:i,j)-1$. Conversely, assume that $S$ is a $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S$ and $x\notin S$ for all $x\in[\min\{k,d^{\prime\prime}\},j)$. So, $|S|=\gamma^{1}_{[1,2]}(\min\{k,d^{\prime\prime}\},j:j)$. Since $x\notin S$ for all $x\in[\min\{k,d^{\prime\prime}\},j)$, $k\leq d$ and the case (i) does not hold, each vertex of $(j,i)$ is adjacent to at most two vertices $i$ and $j$. Hence, $S^{\prime}=S\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}$ and $x\notin S^{\prime}$ for all $x\in[\min\{k,d^{\prime\prime}\},i)\setminus\{j\}$ and so $\gamma^{1}_{[1,2]}(k,i:i,j)\leq|S^{\prime}|=\gamma^{1}_{[1,2]}(\min\{k,d^{\prime\prime}\},j:j)+1$. This completes the proof of the case (ii.b).
@@
    If no vertex of $[d,\min\{k,d^{\prime\prime}\})$ is in $D$, then we get that $D^{\prime}_{1}=D\setminus\{i\}$ is a $[1,2]$-dominating set of $G[1,j]$ such that $j\in D^{\prime}_{1}$ and $x\notin D^{\prime}_{1}$ for all $x\in[d,j)$. Hence,
@@
    If exactly one vertex $w\in[d,\min\{k,d^{\prime\prime}\})$ is in $D$, then we get that $D^{\prime}_{2}=D\setminus\{i\}$ is a $[1,2]$-dominating set of $G[1,j]$ such that $j,w\in D^{\prime}_{2}$ and $x\notin D^{\prime}_{2}$ for all $x\in[d,j)\setminus\{w\}$. Assume that $x^{\prime}\in[d,\min\{k,d^{\prime\prime}\})$. Hence,
@@
    Conversely, assume that $S_{1}$ is a $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j\in S_{1}$ and $x\notin S_{1}$ for all $x\in[d,j)$. So, $|S_{1}|=\gamma^{1}_{[1,2]}(d,j:j)$. Since $x\notin S_{1}$ for all $x\in[d,j)$, $d<k$ and the case (i) does not hold, each vertex of $(j,i)$ is adjacent to at most two vertices $i$ and $j$. Hence, $S^{\prime}_{1}=S_{1}\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}_{1}$ and $x\notin S^{\prime}_{1}$ for all $x\in[k,i)\setminus\{j\}$ and so
@@
    Let $x\in[d,\min\{k,d^{\prime\prime}\})$. Assume that $S_{2}$ is a $[1,2]$-dominating set of $G[1,j]$ with minimum cardinality such that $j,x\in S_{2}$ and $w\notin S_{2}$ for all $w\in[d,j)\setminus\{x\}$. So, $|S_{2}|=\gamma^{1}_{[1,2]}(d,j:j,x)$. We get that $S^{\prime}_{2}=S_{2}\cup\{i\}$ is an $[1,2]$-dominating set of $G[1,i]$ such that $i,j\in S^{\prime}_{2}$ and $w\notin S^{\prime}_{2}$ for all $w\in[k,i)\setminus\{j\}$ and so
@@
Case II. Assume that $j<b$ and $z\geq b$. The proof of cases (iii.a) and (iii.c) is similar to the proof of case (ii.a) and the proof of case (iii.b) is similar to the proof of case (ii.b), respectively.
@@
    Assume that $c<b$. Recall that $k\leq j$, $d=\mathtt{low}(j+1,i-1)$, $d^{\prime}=\mathtt{low}(b,j-1)$ and $k^{\prime}=\min\{k,d,d^{\prime}\}$. Clearly, $k^{\prime}$ exists. Since $b\leq j$, $k^{\prime}\leq b$. We claim that all vertices of $[k^{\prime},j)$ are not in $D$. If $k^{\prime}=k$, the claim is clear. Assume that $k^{\prime}=d$. So, $d\leq b$. Recall that $z=\max\{x\in(j,i):d=\mathtt{low}(x)\}$. We obtain that $z\notin D$ is adjacent to $i,j\in D$. Since $D$ is a $[1,2]$-dominating set, all vertices of $[d,j)$ are not in $D$. Assume that $k^{\prime}=d^{\prime}$. Let $z^{\prime}\in[b,j-1]$ such that $\mathtt{low}(z^{\prime})=d^{\prime}$. Since $c<b$, we get that $z^{\prime}\notin D$ is adjacent to $i,j\in D$ and so all vertices of $[d^{\prime},j)$ are not in $D$. Hence, all vertices of $[k^{\prime},j)$ are not in $D$. The proof of the case (iv.c) is similar to the proof of the case (ii.b).
This completes the proof of the lemma.